September 23, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, September 9, 2015
11:00 am, EDT
JLab: TED 2561B

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at .


  1. Announcements
  2. Summary of Detector Performance
  3. Calibration Data Taking
  4. Subdetector Reports
  5. Simulations
  6. Calibration Processing
  7. AOB


Attending: Sean (NU); Simon, Mark I., Kei, Elton, Justin, Eugene, Alex B., Nathan, Beni, Adesh, Matt S. (JLab); Curtis, Mike S., Will M., Naomi (CMU);

  1. Announcements
    • There was a short meeting yesterday to discussing running this fall. Plans will be more clear by the Collaboration Meeting, but a certain number of standby shifts are being considered. It is not clear how much beam we will get (if any!).
    • The next offline monitoring run will be launched this weekend.
  2. Summary of Detector Performance
    • Curtis said that Matt had sent him some new materials, and the beamline and online groups are working to update their sections.
    • Sean pointed out that efficiencies for several detectors need to be calculated.
  3. Calibration Data Taking
    • Sean presented some preliminary numbers that he collected on how much more data is needed for a full, detailed calibration of all subdetectors. There was substantial discussion.
    • Calorimeter estimates were refined to ~50M pi0's for the BCAL and 10-50M for the FCAL. Sean will follow up about more detailed estimates.
    • About 350M cosmic triggers and the spring data sets were used to study the effects in the CDC, Simon and Mike seemed to think that would be enough. Another shift or so of beam data with magnetic field could be useful.
      • Naomi pointed out that the new fADC125 firmware should be ready on the order of the next couple of weeks. More cosmic data will need to be taken to calibrate the new firmware.
    • Sean will perform some calculations to estimate running times needed.
  4. Subdetector Reports
    • FCAL - Efforts are concentrated on using simulations to understand what factors are driving the detector resolution.
    • BCAL - Elton reported progress on several fronts:
      • George updated the CCDB with effective velocity calibrations and is finalizing the relative time offsets from his analysis.
      • Mark Dalton has updated the gain ratios using the spring data. These ratios are similar to those obtained from the Fall data.
      • Will M. is updating the reconstruction to use single ended BCAL hits. He is currently updating constants, including the new attenuation lengths. The gains he is obtaining so far with the new reconstruction code look similar to those previously obtained.
    • CDC - Simon has been characterizing the "sag" effects in the straw tubes using spring data. He's identified a couple bad connectors, and is working to parameterize the needed corrections. Drift tables with longer drift times may be needed. Mike S. will have more results to report at the tracking meeting tomorrow.
    • FDC - No news.
    • TOF - Nothing new. Sasha is trying to take cosmic data remotely, but is running into trigger problems.
    • Start Counter - Mahmoud is working on the gain calibration using finer bins in z.
    • TAGH - Nathan is finishing up the timewalk corrections using the PSC as a reference. He is finalizing his fits and expects to be finished by the Collaboration Meeting.
    • TAGM - Alex is finishing up his timewalk corrections, which he find to be larger than those in the TAGH. Using just the TAGH to select the correct RF bunch was resulting in picking the wrong bunch a significant number of times, so he is now using the PSC to pick the correct bunch, which is resulting in much cleaner distributions. He is currently refining his results, and will then work on the efficiency determination.
  5. Simulations
    • Mark submitted the first test jobs for the spring
    • Sean will put together files for future simulations together and iterate with Mark.
    • After some short discussion, it was agreed that at the current rate of progress, further updates to mcsmear should wait until HDGeant4 has been validated and released.